OpenStudy (anonymous):

Solve the following equation: -2/3 ln (k-8) + 2/3 ln (k-1) = ln 4

OpenStudy (anonymous):

\[\frac{ -2 }{ 3 }\ln(k-8) + \frac{ 2 }{ 3 }\ln(k-1) = \ln4\]

OpenStudy (anonymous):

Help Anyone?

OpenStudy (anonymous):

can use ... ln x + ln y = ln xy

OpenStudy (anonymous):

Yes I know that but how would you solve that? (k-8)^-2/3 * (k-1)^2/3 = 4

OpenStudy (anonymous):

can factor out 2/3 first...

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

2/3 ln [-(k-8)(k-1)] = ln 4 2/3 ln [(-k+8)(k-1)] = ln 4 \[\sqrt[3]{(-k+8)^2(k-1)^2}=4\]

OpenStudy (anonymous):

Hmmm ok. But then if we're solving it don't we then get a quartic equation

OpenStudy (anonymous):

\[(-k+8)(k-1)=\sqrt{4^3}\]

OpenStudy (anonymous):

\[2/3(\ln(k-1)-\ln(k-8))=\ln4\] \[2/3\ln((k-1)/(k-8))=\ln4\]\[\ln((k-1)/(k-8))^2=\ln64\] \[((k-1)/(k-8))^2=64\] \[(k-1)/(k-8)=8\] \[k-1=8k-64\] \[63=7k\] \[k=9\]

OpenStudy (anonymous):

Oh I see now! I just wasn't too sure about the indice changes. Thanks to both of you!

OpenStudy (anonymous):

np...